OpenStudy (anonymous):

2^x+1+2^x+2^x-1-2^x+2+2^x+3=120

OpenStudy (anonymous):

.... please rewrite.

OpenStudy (anonymous):

2 ^ x +1 +2 ^ x+ 2 ^ x-1-2 ^ x +2 +2 ^ x +3 = 120

OpenStudy (anonymous):

3*2^x +5 = 120 3*2^x = 115 2^x = 115/3 x = log(2)(115/3) x is about 5.26